The partial fractions for


a²/(x(x²+a²)) are

1
Expert's answer
2021-11-26T12:23:15-0500
"\dfrac{a^2}{x(x^2+a^2)}=\dfrac{A}{x}+\dfrac{Bx+C}{x^2+a^2}"

"=\dfrac{Ax^2+Aa^2+Bx^2+Cx}{x(x^2+a^2)}"

"x^2: A+B=0"

"x^1: C=0"

"x^0: Aa^2=a^2"

Then "A=1, B=-1, C=0."

The partial fractions for "\dfrac{a^2}{x(x^2+a^2)}" are



"\dfrac{a^2}{x(x^2+a^2)}=\dfrac{1}{x}-\dfrac{x}{x^2+a^2}"